A mass m, free to move along a line is attracted towards a given point on the line with a force proportional to its distance from the given point . If the mass starts from rest at a distance s from the given. Show that the mass moves in a simple harmonic motion.
1
Expert's answer
2019-03-21T09:58:44-0400

In simple harmonic motion the acceleration of the system, and therefore the net force, is proportional to the displacement and acts in the opposite direction of the displacement.


"F=-kx""F=m\ddot{x}"

Differential equation


"m\ddot{x}+kx=0"

This is a second-order linear differential equation.

The initial conditions: 


"x(0)=s, \dot{x}(0)=0"

Let

"\omega^2=k/m"

The auxiliary equation is


"r^2+\omega^2=0"

Then


"x(t)=A\cos(\omega*t)+B\sin(\omega*t)""x(0)=A=s"

"x(t)=s\cos(\omega*t)+B\sin(\omega*t)"

"\dot{x}(t)=-s\omega\sin(\omega*t)+B\omega\cos(\omega*t)"

"\dot{x}(0)=B\omega=0""B=0"

"x(t)=s\cos(\omega*t)"

"x(t)=s\cos(\sqrt{k/m}*t)"

The equation of a simple harmonic motion.